OpenStudy (anonymous):

anna has a total of 41 coins, which consist of nickels and quaters. her coins hace a total value of $5.45. How many of each coin does she have? set up two equations and then solve

OpenStudy (anonymous):

@tkhunny could you help

OpenStudy (anonymous):

@jim_thompson5910 could you help me

OpenStudy (anonymous):

@Directrix @hartnn can any of you guys help?? im desperate

OpenStudy (anonymous):

x+y=41 0.05x + 0.25y = 5.45 Do Elimination: x+y = 41 4(0.05x+0.25y = 5.45) --> 0.2x+y=21.8 Subtract the like values which is y.: 0.2x-x = 41-21.8 -0.8x = 19.2 Solve for x= 24 Plug x back into original question: x + y = 41 24+y = 41 y=17 x=24 nickels y=17 quarters Check: 25(0.05) +7(0.25) = $5.45

OpenStudy (anonymous):

thanx a ton

OpenStudy (anonymous):

Sure, let me know if you need anymore help.

OpenStudy (anonymous):

i do im about to post my last one